A 200. L kiln is used for vitrifying ceramics. It is currently operating at 1175 ∘C , and the pressure is 1.000 atm . How many moles of air molecules are within the confines of the kiln? Express your answer to three significant figures and include the appropriate units.
Answer #1

T = 1175 + 273.15 = 1448.15 K

n = PV/RT = 1.000atm × 200.L/(0.0821atm-L/K.mol × 1448.15K) = 1.6822mol

Number of moles of air = 1.68 mol. (Answer)